# Towards Better Comprehension of Breaking Changes in the NPM Ecosystem

### v1

@@ Abstract.
Code evolution is prevalent in software ecosystems, which can provide many benefits, such as new features, bug fixes, security patches, etc., while still introducing breaking changes (BCs) that make downstream projects fail to work. Breaking changes cause a lot of effort to both downstream and upstream developers: downstream developers need to adapt to breaking changes and upstream developers are responsible for identifying and documenting them. In the NPM ecosystem, characterized by frequent code changes and a high tolerance for making breaking changes, the effort is larger.
For better comprehension of breaking changes in the NPM ecosystem and to enhance breaking change detection tools, we conduct a large-scale empirical study to investigate breaking changes in the NPM ecosystem. We construct a dataset of explicitly documented breaking changes from 381 popular NPM projects. We find that 93.6% of the detected breaking changes can be covered by developers’ documentation, and about 19% of the breaking changes cannot be detected by regression testing. Then in the process of investigating source code of our collected breaking changes, we yield a taxonomy of JavaScript and TypeScript-specific syntactic breaking changes and a taxonomy of major types of behavioral breaking changes. Additionally, we investigate the reasons why developers make breaking changes in NPM and find three major reasons, i.e., to reduce code redundancy, to improve identifier name, and to improve API design, and each category contains several sub-items.
We provide actionable implications for future research, e.g., automatic naming and renaming techniques should be applied in JavaScript projects to improve identifier names, future research can try to detect more types of behavioral breaking changes. By presenting the implications, we also discuss the weakness of automatic renaming and BC detection approaches, such as the lack of support for public identifiers and various types of breaking changes.
@@ 1. Introduction
The evolution of code is prevalent in software ecosystems (Konstantopoulos et al., 2009; Raemaekers et al., 2012; Lehman, 1980; Traini et al., 2021). Developers of upstream software libraries make code changes to incorporate new features, bug fixes, security patches, component refactorings, and extra-functional improvements (Traini et al., 2021; Tao et al., 2012). However, code evolution may break the contract previously established with its downstream by introducing breaking changes (BCs) in its public APIs, making client applications fail to work (Xavier et al., 2017). For example, renaming frequently used methods or classes can make client projects that depend on the code fail to compile. Therefore, downstream developers will be required to make efforts to adapt to such BCs.
@@
Considering the risks and the effort brought by BCs to downstream developers, many works have been utilized to help developers analyze and measure the impacts of BCs (Mezzetti et al., 2018; Møller and Torp, 2019; Don, [n. d.]). However, they relied on test suites in downstream projects, and either detected JavaScript BCs by directly running test suites (Don, [n. d.]), or via API models [^1] Intuitively, an API model defines the type restrictions of an API (Mezzetti et al., 2018). generated from dynamic execution of test cases (Mezzetti et al., 2018; Møller and Torp, 2019). This leads to the result that BCs observed in prior studies are located in well-known APIs that are popularly used in downstream projects, such as “Client.socket” in socket.io [^2] https://github.com/socketio/socket.io/commit/b73d9be, “Request.prototype.onResponse” in request [^3] https://github.com/request/request/commit/d05b6ba, and each in async. However, BCs are reported to upstream projects because they already manifested themselves in downstream test cases or already caused bugs. For example, in version 3.0.0 of a famous JavaScript project Lodash, the behavior of function “mixin” was changed and not explicitly marked by upstream developers, and a downstream developer reported this issue by self-constructed test code. [^4] https://github.com/lodash/lodash/issues/880 Hence, completely relying on test case execution is not sufficient for understanding BCs.
@@
We observe that in the NPM ecosystem, a number of projects write commit messages complying with Conventional Commits (Con, [n. d.]). According to Conventional Commits, “BREAKING CHANGE” tokens are documented in commit messages to indicate that the corresponding commits contain breaking changes. This is because the core developers of upstream libraries need to ensure the stability of the API and protect the reputation of the upstream projects. When a new code change is submitted for review, they carefully inspect it for the presence of BCs, as well as other bugs or defects. By leveraging the documented BCs provided by developers, there is a potential opportunity to notify downstream users about these BCs.
@@
However, developers experience increased pressure within the NPM ecosystem, primarily driven by frequent code changes and a high tolerance for BCs (Bogart et al., 2021, 2016). Another possible factor is that different developers may have different perceptions of BCs and some of them may ignore the BCs, leading to unaware BCs and causing unforeseen bugs in downstream projects. If we could characterize the large number of documented BCs at source code level, it would assist upstream developers in gaining a deeper understanding of the types of commits that could potentially introduce BCs. Furthermore, such categorization would provide valuable insights for researchers to develop tools aimed at detecting BCs in the future.
@@
To bridge this gap, we conduct an empirical study to better comprehend documented BCs. We first select 381 popular JavaScript projects (randomly selected from 35,786 projects with more than 50 GitHub stars in the NPM ecosystem) and clone their associated repositories. We then identify BC-related commits admitted by developers by searching for BREAKING CHANGE tokens in commit messages. For the 5,242 identified BC commits, we manually select 1,519 commits that 1) containing not too long commit messages, 2) related to JavaScript production code and 3) associated with documentation that can explain the reasons behind the commit. These selected BC commits are distributed across 131 distinct projects and the projects can be grouped into six categories according to functionalities and usages, i.e., utilities, frontend projects, development-related tools, database-related tools, plugins and Web API related tools (detailed in Section 3), which makes the BC commits in our projects representative. To extract BCs from these BC commits, we build BC type categories by learning from a previous API detection tool for Java (Brito et al., 2018) and adding new JavaScript and TypeScript-specific features. Our taxonomy includes Remove, Rename, Change Signature, Change Behavior types, etc., which is detailed in Section 2.2. Through several thematic analyses, we obtain JavaScript and TypeScript specific code features (RQ2 and RQ3) and a taxonomy of reasons behind BCs. We answer the following research questions in this study:
@@
RQ1: To what extent do detected breaking changes and documented breaking changes overlap? In our work, we collect BCs from developers’ intention, i.e., the explicitly documented BCs. We first check whether documented BCs actually break test code. Therefore, we do regression testing on our collected BC commits, similar to previous works (Mezzetti et al., 2018; Møller and Torp, 2019; Mostafa et al., 2017; Zhang et al., 2022). We find 97% of detected BCs are documented and 81% of the documented BCs can be detected by regression testing. The result shows that most detected BCs are well documented, while a proportion of documented BCs cannot be detected by regression testing. Since the documented BCs cover most of the detected BCs, hence extracting BCs from documentation (especially commit messages, issues, and pull requests) is reasonable.
@@
RQ2: What syntactic breaking changes in the NPM ecosystem are specific to JavaScript and TypeScript? We investigate whether there are some BCs related to JavaScript-specific language features. Despite the triviality of most syntactic BCs (such as moving classes to another package, removing a field of a class), we find some JavaScript-specific BCs are notable, including two removing operations, i.e., 1) removing default export, 2) removing export of an element in one place, and five types of JavaScript and TypeScript specific signature changes, e.g., parameter changes in configuration objects, switch between callback and Promise, etc.
@@
RQ3: How do developers make behavioral breaking changes? In this RQ, we investigate how developers perform BCs at source code level. Since syntactic BCs can be detected through syntax analysis and code refactoring detection, we mainly focus on behavioral BCs (a.k.a, semantic BCs in some works) that change the internal program logic. We find four major types of Change Behavior BC, i.e., 1) changing the specification return values, 2) changing process for some option values, 3) changing default or initial value of variables and 4) changing error handling method.
@@
RQ4: Why do developers make breaking changes in NPM ecosystem? In this RQ, we revisit the rationale of BCs in the NPM ecosystem and find three main factors (each with some sub-factors) that motivate developers to make BCs, i.e., 1) to reduce code redundancy, 2) to improve identifier name, 3) to improve API design, and divide each reason into some sub-items, which extended the previous works (Bogart et al., 2016, 2021; Brito et al., 2020).
@@
Based on our empirical findings, we provide actionable implications for future research, including 1) automatic renaming deserves much concern in JavaScript projects since poor identifier name is a main contributor to technical debt, 2) automatic tools for ensuring code consistency are necessary, 3) future research can try to detect more types of behavioral BCs. By presenting these implications, we also discuss the weakness of automatic renaming and BC detection approaches, such as lacking support for public identifiers and various types of BCs.
@@
    We empirically identify how developers perform BCs at source code level, yielding notable findings like JavaScript and TypeScript specific syntactic BCs and typical actions of behavioral BCs. We also investigate why developers perform BCs in the NPM ecosystem.
@@
Data availability. We provide the datasets of our research at https://zenodo.org/records/13219136.
@@ 2.1. Preliminary of Breaking Changes
Prior to our work, a number of studies have uncovered BCs from many aspects, such as the occurrence of BCs, the impacts of BCs on downstream projects, and motivations for making BCs. The research works (Decan and Mens, 2019; Venturini et al., 2023; Raemaekers et al., 2017) have demonstrated the widespread occurrence of BCs in the NPM ecosystem and their effects on client applications. Venturini et al. found that in their sampled packages, 11.7% of all client packages and 13.9% of their releases are impacted by breaking changes, and notably, 44% of the breaking changes are introduced in minor or patch releases (according to Semantic Versioning (Sem, [n. d.]), developers should not introduce breaking changes in non-major releases) (Venturini et al., 2023). Bogart et al. found that coarse-grained motivations for making BCs include requirements and context changes, bugs and new features, rippling effects from upstream changes, and technical debt from postponed changes (Bogart et al., 2016, 2021).
@@
However, despite much knowledge of BC, to the best of our knowledge, none of the studies precisely defined BC. Researchers collected BCs in various forms in previous studies. For example, in the evaluation of several breaking change detection tools (Zhang et al., 2022; Mezzetti et al., 2018), researchers collected BCs by running downstream test cases: if a test case of a downstream project could not pass with a newer provider (after the code change), then the code change was identified as a BC. However, on the one hand, not all providers are dependent on many client applications. If a code change has no triggering test case in client applications, we cannot determine whether it is actually incompatible. On the other hand, client code may not follow the specification of providers and incorrectly access APIs, e.g., passing an improper value to a function, which will result in undefined behavior (Møller and Torp, 2019), including test failure, but it does not reflect a BC. In another breaking change detection tool APIDiff (Brito et al., 2018), Brito et al. utilized code refactoring detection tool RefDiff (Silva and Valente, 2017) to identify syntax-related BCs (e.g., changing method signature or renaming class), however, as they pointed out, the detected BCs are just breaking change candidates (BCC), since some syntax changes are applied to internal methods and classes, which are not intended for public use. To this end, they also asked developers to check whether the detected BCCs are actually BCs (Brito et al., 2020).
@@
In our study, we adopt Brito et al.’s criteria (Brito et al., 2020), since we also investigate the BCs from developers’ perspective, i.e., how and why they perform BCs. Specifically, a BC should 1) be confirmed by developers, 2) be categorized into pre-defined BC types. To achieve the first criterion, we identify BCs by searching explicit BC declarations in the documentation (usually commit messages). To achieve the second criterion, we define several BC types for JavaScript and TypeScript on the basis of Brito et al.’s Java BC types (Brito et al., 2018) and JavaScript development experience. We detail the BC types in Section 2.2 and describe how to extract BCs in Section 3.3.
@@ 2.2. Types of Breaking Changes in NPM Projects
Table 1. Syntactic Breaking Change Types Adopted from Previous Works
@@
Existing BC detection tools for Java, typically APIDiff (Brito et al., 2018), support many syntactic and object-oriented programming (OOP) related BC types, such as remove classes and push down fields. Since OOP features are also supported in JavaScript (since ECMAScript 2015 (ECM, [n. d.]), class can be directly used while legacy code must leverage prototype mechanism (inh, [n. d.])), we learn from the BC types for Java language and make them adapt to JavaScript and TypeScript (a superset of JavaScript, in practice many source code files are written in TypeScript and will be compiled into JavaScript files) by retaining available BC actions in JavaScript and adding more BCs which are not considered previously. We then construct a category of JavaScript and TypeScript BCs (shown in Table 1), where interface, enum and type are only available in TypeScript. We explain each type as follows:
@@
    Rename BC refers to identifier changes of public code elements. Rename BCs can be applied to almost all code elements, such as exported classes, interfaces, and enums, as well as methods and fields in exported classes and interfaces.
@@
    Remove BC refers to removal of public code elements. Remove BC can be applied to almost all code elements. After a Remove BC, the affected public element is not accessible by downstream applications, e.g., an exported class is completely removed, or no longer marked as exported.
@@
    Move BC can be applied to all APIs supported by Rename BCs. For example, a class is moved to another package since its functionality is more related to that package.
@@
    Inline is a refactoring operation in object-oriented programming: removing a public method and copying its body into an existing method. Simple methods are often inlined to make code more straightforward. In fact, Inline can be regarded as a special case of Remove, since the inlined public method is removed from public access. By contrast, normal “Remove” actions do not copy the content of the removed method to another existing method. However, since inline is a very common refactoring operation in Object-oriented languages, we still regard it as a separate type of BC, like previous works (Brito et al., 2018; Silva and Valente, 2017).
@@
    Change Signature BC refers to the modification of method signatures except for directly changing method names. Adding or removing modifiers (such as “static”, “private”) and changing the parameters are typical refactoring actions of this type.
Besides syntactic BCs, Change Behavior BC (also called semantic BC in some works (Mezzetti et al., 2018; Zhang et al., 2022)) refers to those operations changing the internal behavior (e.g., programmatic logic in methods, and content of global constants that can be used by some methods) rather than syntactic elements, such as changing class names and adding required parameters. In other words, after a behavioral BC, the affected classes (or methods, interfaces, etc.) can be invoked as the same way as before.
In our work, we focus on BCs in JavaScript source code, hence we do not take BCs in non-source code files into consideration, typically CSS, Markdown, HTML, package.json files and configuration files related to TypeScript, Webpack and ESLint, etc.

### v2

@@ Abstract.
Code evolution is prevalent in software ecosystems, which can provide many benefits, such as new features, bug fixes, security patches, etc., while still introducing breaking changes that make downstream projects fail to work. Breaking changes cause a lot of effort to both downstream and upstream developers: downstream developers need to adapt to breaking changes and upstream developers are responsible for identifying and documenting them. In the NPM ecosystem, characterized by frequent code changes and a high tolerance for making breaking changes, the effort is larger.
For better comprehension of breaking changes in the NPM ecosystem and to enhance breaking change detection tools, we conduct a large-scale empirical study to investigate breaking changes in the NPM ecosystem. We construct a dataset of explicitly documented breaking changes from 381 popular NPM projects. We find that 95.4% of the detected breaking changes can be covered by developers’ documentation, and 19% of the breaking changes cannot be detected by regression testing. Then in the process of investigating source code of our collected breaking changes, we yield a taxonomy of JavaScript and TypeScript-specific syntactic breaking changes and a taxonomy of major types of behavioral breaking changes. Additionally, we investigate the reasons why developers make breaking changes in NPM and find three major reasons, i.e., to reduce code redundancy, to improve identifier names, and to improve API design, and each category contains several sub-items.
We provide actionable implications for future research, e.g., automatic naming and renaming techniques should be applied in JavaScript projects to improve identifier names, future research can try to detect more types of behavioral breaking changes. By presenting the implications, we also discuss the weakness of automatic renaming and breaking change detection approaches, such as the lack of support for public identifiers and various types of breaking changes.
@@ 1. Introduction
The evolution of code is prevalent in software ecosystems (Konstantopoulos et al., 2009; Raemaekers et al., 2012; Lehman, 1980; Traini et al., 2021). Developers of upstream software libraries make code changes to incorporate new features, bug fixes, security patches, component refactorings, and extra-functional improvements (Traini et al., 2021; Tao et al., 2012). However, code evolution may break the contract previously established with its downstream by introducing breaking changes (BCs) in its public APIs, making client applications fail to work (Xavier et al., 2017). For example, renaming frequently used methods or classes can make client projects that depend on the code fail to compile. Therefore, downstream developers will be required to make efforts to adapt to such breaking changes.
@@
Considering the risks and the effort brought by breaking changes to downstream developers, many works have been utilized to help developers analyze and measure the impacts of breaking changes (Mezzetti et al., 2018; Møller and Torp, 2019; Don, [n. d.]). However, they relied on test suites in downstream projects, and either detected JavaScript breaking changes by directly running test suites (Don, [n. d.]), or via API models [^1] Intuitively, an API model defines the type restrictions of an API (Mezzetti et al., 2018). generated from dynamic execution of test cases (Mezzetti et al., 2018; Møller and Torp, 2019). This leads to the result that breaking changes observed in prior studies are located in well-known APIs that are popularly used in downstream projects, such as Client.socket in socket.io [^2] https://github.com/socketio/socket.io/commit/b73d9be, Request.prototype.onResponse in request [^3] https://github.com/request/request/commit/d05b6ba, and each in async. However, breaking changes are reported to upstream projects because they already manifested themselves in downstream test cases or already caused bugs. For example, in version 3.0.0 of a famous JavaScript library lodash, the behavior of function mixin was changed and not explicitly marked by upstream developers, and a downstream developer reported this issue by self-constructed test code. [^4] https://github.com/lodash/lodash/issues/880 Hence, completely relying on test case execution is not sufficient for understanding breaking changes.
@@
We observe that in the NPM ecosystem, a number of projects write commit messages complying with Conventional Commits (Con, [n. d.]). According to Conventional Commits, BREAKING CHANGE tokens are documented in commit messages to indicate that the corresponding commits contain breaking changes. This is because the core developers of upstream libraries need to ensure the stability of the API and protect the reputation of the upstream projects. When a new code change is submitted for review, they carefully inspect it for the presence of breaking changes, as well as other bugs or defects. By leveraging the documented breaking changes provided by developers, there is a potential opportunity to notify downstream users about these breaking changes.
@@
However, developers experience increased pressure within the NPM ecosystem, primarily driven by frequent code changes and a high tolerance for breaking changes (Bogart et al., 2021, 2016). Another possible factor is that different developers may have different perceptions of breaking changes and some of them may ignore the breaking changes, leading to unaware breaking changes that cause unforeseen bugs in downstream projects. If we could characterize the large number of documented breaking changes at source code level, it would assist upstream developers in gaining a deeper understanding of the types of commits that could potentially introduce breaking changes. Furthermore, such categorization would provide valuable insights for researchers to develop tools aimed at detecting breaking changes in the future.
@@
To bridge this gap, we conduct an empirical study to better comprehend documented breaking changes. We first select popular projects with over 50 GitHub stars and associated links to an existing GitHub repository, resulting in a total of 35,786 repositories. We then randomly select 381 repositories (95% confidence level and 5% margin of error) for further investigation. We then clone their associated repositories. After this, we identify BC-related commits admitted by developers by searching for BREAKING CHANGE tokens in commit messages. For the 5,242 identified BC commits, we manually select 1,519 commits that 1) do not contain too long commit messages, 2) are related to JavaScript production code and 3) are associated with documentation that can explain the reasons behind the commit. These selected BC commits are distributed across 131 distinct projects and the projects can be grouped into six categories according to functionalities and usages, i.e., utilities, frontend projects, development-related tools, database-related tools, plugins and Web API related tools (detailed in Section 3), which makes the BC-related commits in our projects representative. To extract breaking changes from these BC commits, we build breaking change type taxonomy by learning from a previous API detection tool for Java (Brito et al., 2018) and adding new JavaScript and TypeScript-specific features. Our taxonomy includes Remove, Rename, Change Signature, Change Behavior types, etc., which is detailed in Section 2.2. Through several thematic analyses, we obtain JavaScript and TypeScript specific code features (RQ2 and RQ3) and a taxonomy of reasons behind breaking changes. We answer the following research questions in this study:
@@
RQ1: To what extent do detected breaking changes and documented breaking changes overlap? In our work, we collect breaking changes from developers’ intention, i.e., the explicitly documented breaking changes. We first check whether documented breaking changes actually break test code. Therefore, we do regression testing on our collected BC-related commits, similar to previous works (Mezzetti et al., 2018; Møller and Torp, 2019; Mostafa et al., 2017; Zhang et al., 2022). We find 95.4% of detected breaking changes are documented and 81% of the documented breaking changes can be detected by regression testing. The result shows that most detected breaking changes are well documented, while a proportion of documented breaking changes cannot be detected by regression testing. Since the documented breaking changes cover most of the detected breaking changes, it is reasonable to extract breaking changes from documentation (especially commit messages, issues, and pull requests).
@@
RQ2: What syntactic breaking changes in the NPM ecosystem are specific to JavaScript and TypeScript? We investigate whether there are some breaking changes related to JavaScript-specific language features. Despite the triviality of most syntactic breaking changes (such as moving classes to another package, removing a field of a class), we find some JavaScript-specific BCs are notable, including two removing operations, i.e., 1) removing default export, 2) removing export of an element in one place, and five types of JavaScript and TypeScript specific signature changes, e.g., parameter changes in configuration objects, switch between callback and Promise, etc.
@@
RQ3: How do developers make behavioral breaking changes? In this RQ, we investigate how developers perform breaking changes at source code level. Since syntactic breaking changes can be detected through syntax analysis and code refactoring detection, we mainly focus on behavioral breaking changes (a.k.a, semantic breaking changes in some works) that change the internal program logic. We find four major types of Change Behavior breaking changes, i.e., 1) changing the specification return values, 2) changing process for some option values, 3) changing default or initial value of variables and 4) changing error handling method.
@@
RQ4: Why do developers make breaking changes in NPM ecosystem? In this RQ, we revisit the rationale of breaking changes in the NPM ecosystem and find three main factors (each with some sub-factors) that motivate developers to make breaking changes, i.e., 1) to reduce code redundancy, 2) to improve identifier names, 3) to improve API design, and divide each reason into some sub-items, which extends the previous works (Bogart et al., 2016, 2021; Brito et al., 2020).
@@
Based on our empirical findings, we provide actionable implications for future research, including 1) automatic renaming deserves much concern in JavaScript projects since poor identifier naming is a main contributor to technical debt, 2) automatic tools for ensuring code consistency are necessary, 3) future research should strive to detect more types of behavioral breaking changes. By presenting these implications, we also discuss the weakness of automatic renaming and breaking change detection approaches, such as lacking support for public identifiers and various types of breaking changes.
@@
    We empirically identify how developers perform breaking changes at source code level, yielding notable findings like JavaScript and TypeScript specific syntactic breaking changes and typical actions of behavioral breaking changes. We also investigate why developers perform breaking changes in the NPM ecosystem.
@@
Data availability. We provide the replication package of our research at https://doi.org/10.5281/zenodo.13927690.
@@ 2.1. Preliminary of Breaking Changes
Prior to our work, a number of studies have uncovered breaking changes from many aspects, such as the occurrence of breaking changes, the impacts of breaking changes on downstream projects, and motivations for making breaking changes. The research works (Decan and Mens, 2019; Venturini et al., 2023; Raemaekers et al., 2017) have demonstrated the widespread occurrence of breaking changes in the NPM ecosystem and their effects on client applications. Venturini et al. found that in their sampled packages, 11.7% of all client packages and 13.9% of their releases are impacted by breaking changes, and notably, 44% of the breaking changes are introduced in minor or patch releases (according to Semantic Versioning (Sem, [n. d.]), developers should not introduce breaking changes in non-major releases) (Venturini et al., 2023). Bogart et al. found that coarse-grained motivations for making breaking changes include requirements and context changes, bugs and new features, rippling effects from upstream changes, and technical debt from postponed changes (Bogart et al., 2016, 2021).
@@
However, despite much knowledge of BC, to the best of our knowledge, none of the studies precisely defined BC. Researchers collected breaking changes in various forms in previous studies. For example, in the evaluation of several breaking change detection tools (Zhang et al., 2022; Mezzetti et al., 2018), the authors collected breaking changes by running downstream test cases: if a test case of a downstream project could not pass with a newer provider (after the code change), then the code change was identified as a BC. However, on the one hand, not all providers are dependent on many client applications. If a code change has no triggering test case in client applications, we cannot determine whether it is actually incompatible. On the other hand, client code may not follow the specification of providers and incorrectly access APIs, e.g., passing an improper value to a function, which will result in undefined behavior (Møller and Torp, 2019), including test failure, but it does not reflect a BC. In another breaking change detection tool APIDiff (Brito et al., 2018), Brito et al. utilized the code refactoring detection tool RefDiff (Silva and Valente, 2017) to identify syntax-related breaking changes (e.g., changing method signature or renaming class). However, as they pointed out, the detected breaking changes are just breaking change candidates (BCC), since some syntax changes are applied to internal methods and classes, which are not intended for public use. To this end, they also asked developers to check whether the detected BCCs are actually breaking changes (Brito et al., 2020).
@@
In our study, we adopt Brito et al.’s criteria (Brito et al., 2020), since we also investigate the breaking changes from developers’ perspective, i.e., how and why they perform breaking changes. Specifically, a breaking change should 1) be confirmed by developers, 2) be categorized into pre-defined BC types. To achieve the first criterion, we identify breaking changes by searching explicit BC declarations in the documentation (usually commit messages). To achieve the second criterion, we define several breaking change types for JavaScript and TypeScript on the basis of Brito et al.’s Java BC types (Brito et al., 2018) and JavaScript development experience. We detail the BC types in Section 2.2 and describe how to extract breaking changes in Section 3.3.
@@ 2.2. Types of Breaking Changes in NPM Projects
Table 1. Breaking Change Types Adopted from Previous Works
@@
Change Behavior the content of method, field, constant and variable
Existing BC detection tools for Java, typically APIDiff (Brito et al., 2018), support many syntactic and object-oriented programming (OOP) related BC types, such as remove classes and push down fields. Since OOP features are also supported in JavaScript (since ECMAScript 2015 (ECM, [n. d.]), class can be directly used while legacy code must leverage the prototype mechanism (inh, [n. d.])), we learn from the breaking change types for the Java language and adapt them to JavaScript and TypeScript (a superset of JavaScript, in practice many source code files are written in TypeScript and will be compiled into JavaScript files) by retaining available BC actions in JavaScript and adding more breaking change types which are not considered previously. We then construct a taxonomy of JavaScript and TypeScript breaking changes (shown in Section 1), where interface, enum and type are only available in TypeScript. We explain each type as follows:
@@
    Rename refers to identifier changes of public code elements. Rename BCs can be applied to many code elements, such as exported classes, interfaces, and enums, as well as methods and fields in exported classes and interfaces.
@@
    Remove refers to removal of public code elements. Remove BC can be applied to many code elements. After a Remove BC, the affected public element is not accessible by downstream applications, e.g., an exported class is completely removed, or no longer marked as exported.
@@
    Move can be applied to all APIs supported by Rename BCs. For example, a class is moved to another package since its functionality is more related to that package.
@@
    Inline is a refactoring operation in object-oriented programming: removing a public method and copying its body into an existing method. Simple methods are often inlined to make code more straightforward. In fact, Inline can be regarded as a special case of Remove, since the inlined public method is removed from public access. By contrast, normal “Remove” actions do not copy the content of the removed method to another existing method. However, since inline is a very common refactoring operation in object-oriented languages, we still regard it as a separate breaking change type, like previous works (Brito et al., 2018; Silva and Valente, 2017).
@@
    Change Signature refers to the modification of method signatures except for directly changing method names. Adding or removing modifiers (such as “static”, “private”) and changing the parameters are typical refactoring actions of this type.
  7. (7)
    Change Behavior (also called semantic breaking changes in some works (Mezzetti et al., 2018; Zhang et al., 2022)) refers to those operations changing the internal behavior (e.g., programmatic logic in methods, and content of global constants that can be used by some methods) rather than syntactic elements, such as changing class names and adding required parameters. In other words, after a behavioral breaking change is performed, the affected classes (or methods, interfaces, etc.) can be invoked as the same way as before.
In our work, we focus on breaking changes in JavaScript source code, hence we do not take breaking changes in non-source code files into consideration, typically CSS, Markdown, HTML, package.json files and configuration files related to TypeScript, Webpack and ESLint, etc.